OpenStudy (anonymous):

H12P3: Our specification is that we want there to be two break frequencies for this filter: 10Hz and 15kHz. We also want midband frequencies, such as signals around 1kHz, to be amplified by about a factor of 2. We have available a 1.5uF and a 470pF capacitor. We also have a variety of resistors between 100 Ohms and 100 kOhms . However, our op amp really likes RS = 20 kOhms . Your job is to fill in the remaining details of this design.

OpenStudy (anonymous):

part1) First find C1=1.5u(answer of part 4) Then R1*C1=1/(2*pi*f) R1*C1=1/(2*pi*10) so R1= 10610 part 2)First find C2=470p (answer of part 5) R2*C2=1/(2*pi*15*10^3) R2=22575 part 3) Rs = Rf = 20kΩ = 20000 Ω part 6 : R3= R1 part 7 : R4= R2 part 8: c3=c1 part 9 : c4= c2 part 10 :Midband gain = ­-(R4/R3)=-2.1277

OpenStudy (anonymous):

Thank you so much WAK, for your help in making it clear :)

OpenStudy (anonymous):

welcome :)